OpenStudy (anonymous):

evaluate 48 * 2 to the power of -3 free medals

OpenStudy (anonymous):

Can you draw the equation?

OpenStudy (anonymous):

I'm guessing \(\large 48 \cdot 2^{-3}\)

OpenStudy (anonymous):

48\[48 * 2^{-3}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

What does it mean to have a negative exponent, and what is 2^3 anyway?

OpenStudy (anonymous):

1/8

OpenStudy (anonymous):

is it 1/8

OpenStudy (anonymous):

the answer is 6

OpenStudy (anonymous):

oh ok thx

OpenStudy (anonymous):

Can you give a more detailed explanation please, @anas2000 , just to make sure @cool987654321 understands?

OpenStudy (anonymous):

okay but how do you explain all i got was 48*1/8=6

OpenStudy (anonymous):

That's good enough, @anas2000 . Just showing that it is equivalent to 48 ÷ 8 = 6 is important for understanding.

OpenStudy (anonymous):

okay